CRY+VIS Interface Meeting
Date: 2022/3/25 13:00-13:40
Participants: Takahashi, Yokozawa, Hirata, Ikeda, Miyoki, Aso, Sato, Tomaru, Ushiba, Yasui, Kimura, Uchiyama
- IY: Height adjustment. H2 geophone. TM actuator.
- EY: Recovery of GAS filters.
- Type-B/Bp/C: Health check after peeling of FC in SR2. Leakage at the feedthrough flange in IMC, replacement was scheduled on 4/7.
- Replacement of flip cables: TMS.
- TM height setting: ITMY.
- Oplev: Summarized the remain works.
- Commissioning in air: Alignmenet of PR2. Alignment of IMC input beam.
- Commissioning in vacuum:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: Leak check in EY. Opening the GV in EXV.
- Heater control: Test in IXV/EXV.
Date: 2022/3/25 14:00-14:15
Participants: Takahashi, Terrence, Ushiba, Aso, Tanaka
- Recovered the GAS filters in EY.
- Investigated the H2 geophone in IY.
- Adjusted the TM height in IY.
- Health check after peeling of FC in SR2.
- Lekage at the feedthrough flange was found in IMC. The flange will be replaced on 4/7.
- BIO switches for the IP (H3) dewhitening filter was broken in IY.
- Type-A team:
- Type-B team:
(Date in Kamioka)
The week from 4/4
- Takahashi: 4/5-8
- Sato: 4/5-8
- Hirata: 4/5-8
The week from 4/11
- Takahashi: 4/12-15
- Sato: 4/12-15
- Hirata: 4/12-15